OpenStudy (anonymous):

Calculate the ΔG for the following system. Then state if the system is spontaneous or not spontaneous. ΔH = + 25 kJ ΔS = +5.0 J/K T = 23°C 1.5 x 103 kJ; not spontaneous +24 kJ; not spontaneous +25 kJ; not spontaneous 1.5 x 103 kJ; spontaneous +24 kJ; spontaneous +25 kJ; spontaneous

OpenStudy (nurali):

Note that T = 273 + 23 = 296 K. Then, by Gibb's Free Energy: ΔG = ΔH - TΔS = (+25 kJ) - (296 K)(+5.0 J/K) = (+25 kJ) - (296 K)(+0.005 kJ/K), by converting ΔS to kJ/K = 23.52 kJ > 0. Since ΔG > 0, the reaction is not spontaneous.

OpenStudy (anonymous):

i still dont get it

OpenStudy (nurali):

+24 kJ; not spontaneous

OpenStudy (anonymous):

thank u

OpenStudy (nurali):

My Pleasure.